OpenStudy (anonymous):

What value of s makes the statement true? (–7 + s) + (–14) = (12 + (–7)) + (–14) A. –30 B. –12 C. –7 D. 12

OpenStudy (igreen):

@blake57roger Don't give out direct answers.

OpenStudy (anonymous):

ha its 12 its just switched arround

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

did i say that out loud?

OpenStudy (igreen):

Plug in 12: (–7 + 12) + (–14) = (12 + (–7)) + (–14) Add: 5 - 14 = 12 - 7 - 14 Subtract: -9 = 12 - 7 - 14 Subtract: -9 = 5 - 14 Can you subtract 5 - 14? @christine_

OpenStudy (anonymous):

9

OpenStudy (igreen):

No, it's -9. So that gives you -9 = -9, which is true..so 12 is the answer.

OpenStudy (igreen):

@christine_